# Weighted Birkhoff averages: Deterministic and probabilistic perspectives

### v2

@@ 1.1.1 The weighted quasi-Monte Carlo method
$f=\sum\limits_{k\in\mathbb{Z}_{*}^{\infty}}{\hat{f}(k)\exp\left({2\pi i\left\langle{k,x}\right\rangle}\right)}\quad\text{with}\quad\|f\|_{\sigma,\infty}:=\sum\limits_{k\in\mathbb{Z}_{*}^{\infty}}{\|{\hat{f}(k)}\|_{\ell^{\infty}}{\exp\left(2\pi\sigma{{|k|_{\eta}}}\right)}}<+\infty.$
@@
    The discrete case: for any $0\neq k\in\mathbb{Z}_{*}^{\infty}$ and $n\in\mathbb{Z}$, $\left|{\left\langle{k,\rho}\right\rangle-n}\right|>\gamma{\left({\prod\nolimits_{j\in{\mathbb{N}}}{(1+{{\left\langle j\right\rangle}^{\tau}}{{\left|{{k_{j}}}\right|}^{\tau}})}}\right)^{-1}}$, provided $\tau>1$;
@@
In contrast to the delicate results in finite dimensions, it seems that there is no known so-called “0–1 law” for nonresonance in infinite dimensions. However, the infinite-dimensional Diophantine nonresonance in Definition 1.4 is universal. More precisely, under the standard probability measure $\mathbb{P}$ induced by the product measure of the cube $\left[{1,2}\right]^{\mathbb{\infty}}$} (equivalent to $\mathbb{T}^{\infty}$}), the set of these rotations admits full probability measure [Bou05, MP21]. This is sufficient for us to establish the universal exponential convergence of weighted Birkhoff averages, via analytic observables. It should be emphasized that the universal nonresonance is not necessarily confined to the aforementioned Diophantine form. One can construct more general cases, which we do not pursue here.
@@
      If the observable $f$ is “sub-super analytic” on $\mathbb{T}^{d}$, namely $\|\hat{f}(k)\|_{\ell^{\infty}}=\mathcal{O}\left(\exp(-\|k\|_{\ell^{1}}^{v})\right)$ for any $v>1$, then for any given $0<\kappa<1$, when $p,q$ are sufficiently large, for almost all rotations $\rho\in\mathbb{R}^{d}$, the convergence can be of arbitrary exponential type. More precisely, the error is $\mathcal{O}\left({\exp({-{N^{\kappa}}})}\right)$.
@@
      If the observable $f$ is “super analytic” on $\mathbb{T}^{\infty}$, namely $\|\hat{f}(k)\|_{\ell^{\infty}}=\mathcal{O}\left(\exp(-\exp(|k|_{\eta})\right)$, then for almost all rotations $\rho\in\mathbb{R}^{\infty}$, the convergence can be of exponential type. More precisely, the error is $\mathcal{O}\left({\exp({-{N^{\zeta_{4}}}})}\right)$ for any $0<{\zeta_{4}}<{\left({1+1/\min\left\{{p,q}\right\}}\right)^{-1}}$.
@@
      holds for some universal constant $c>0$, where ${\zeta_{5}}={\left({1+1/\min\left\{{p,q}\right\}}\right)^{-1}}$. Moreover, the universal and implied constants only depend on $T$ and ${\max_{0\leqslant n\leqslant T}}\left|{f\left(n\right)}\right|$.
@@
As previously emphasized, although the results here are stated specifically for the case of toral translation, they can be readily formulated in a conjugate manner [DSSY17, DY18, DM23, RB24] to extend to general nonlinear dynamical systems (for example, those that can be smoothly conjugated to toral translation through KAM theory, etc.). Consequently, Theorem 1.2 has broad applicability, particularly in celestial mechanics, since the motion of celestial bodies is often quasi-periodic. For the sake of brevity, we prefer not to complicate the issue further.
@@
The terms “quasi-periodic” and “almost periodic” are used here because the corresponding rotations are nonresonant; for example, they satisfy certain finite-dimensional or infinite-dimensional Diophantine conditions as in Definitions 1.2 and 1.4. To simplify the exposition, we prefer not to state these concepts in Theorem 1.2. These rotations are universal in the sense of measure, meaning that the set of them admits full Lebesgue measure or full probability measure, as previously mentioned. This weighted quasi-Monte Carlo method is not only effective for quasi-periodic, almost periodic, and periodic cases, but also yields similar results for general decaying waves [TL25a, Theorems 1.1 and 1.2]. As a corollary, when the observational data can be decomposed into a linear combination of quasi-periodic, almost periodic, and periodic components, as well as decaying waves, this method also exhibits excellent acceleration effects. Here, we also mention that the “super-analyticity” imposed for the quasi-periodic case in (II)–(6) of Theorem 1.2 is not as strong as it might appear at first glance (of course, the corresponding “super-analyticity” in the finite-dimensional case is indeed strong). This is because, unlike the finite-dimensional case, the infinite-dimensional case requires consideration of the cardinality estimates brought by the $\mathbb{Z}_{*}^{\infty}$ lattice, which inevitably leads to stronger decay of individual Fourier coefficients.
@@
To facilitate the reader’s understanding of Theorem 1.2, we present the numerical simulations in Figure 4. For convenience, we denote the actual error in the $\ell^{\infty}$-norm as ${\rm{Error}}^{p,q}_{N}$. We consider a quasi-periodic orbit generated by $f(x)=\sin(2\pi x)+\cos(10\pi x)$, with an initial value $\theta=0$ and a rotation number $\rho=1/2\pi$, which is universal (i.e., not specifically constructed to achieve the “best” convergence). The choice of $f$ as a finite-order trigonometric polynomial is motivated by practical considerations: in numerical simulations, we cannot truly achieve infinity, and extremely high orders would lead to inaccurate computations. According to Theorem 1.2, specifically the part (I)–(4), the theoretical error is $\mathcal{O}\left(\exp(-cN^{\zeta_{2}})\right)$, where ${\zeta_{2}}={\left({1+1/\min\left\{{p,q}\right\}}\right)^{-1}}$ and $c>0$ is a universal constant that may depend on $p$ and $q$. To visualize the results more intuitively, we take the logarithm of both the actual error and the theoretical error (and thus the theoretical error becomes $\mathcal{O}\left(-N^{\zeta_{2}}\right)$, where ${\zeta_{2}}={\left({1+1/\min\left\{{p,q}\right\}}\right)^{-1}}$). This transformation results in curves that are more curved than $\mathcal{O}^{\#}(-N)$. We present four cases: $p=q=1/2$, $p=q=1$, $p=1$ and $q=2$, and $p=q=2$. It can be observed from Figure 4 that as $\min\{p,q\}$ increases, the exponential convergence rate of the actual error indeed accelerates as we proved, closely matching the changes in the order of the theoretical error (although we do not estimate the control constant $c>0$ as in [TL25a] for general decaying waves, which would have provided a more precise result). Notably, for the cases $p=q=1$ and $p=1$, $q=2$, the actual errors are almost identical, aligning well with our theoretical conclusions, as $\min\{1,1\}=\min\{1,2\}=1$ at this point.
In addition to the results in the general sense mentioned above, [TL24a, Theorems 2.1, 2.2, 3.1 and 3.2] and [TL24c, Theorems 3.1–3.4] also investigate results based on Liouville rotations. Specifically, for any Liouville type rotation, as long as the observable is sufficiently smooth, the weighted Birkhoff average can converge at an exponential rate. This is quite different from the phenomena observed in classical dynamical systems: Firstly, many dynamical systems essentially impose restrictions on the irrationality of rotations, such as the Bruno condition, and do not allow arbitrary Liouville rotations. Secondly, even for the specific case of irrational translations, without weighting, such results are absolutely impossible according to the co-homological equation argument. However, when the regularity of the observable is insufficient, such weak nonresonance will reduce the convergence rate in the weighted case, which is entirely reasonable from the perspective of dynamical systems.
@@
There exist weighting functions $w^{*}(x)\in C_{0}^{\infty}([0,1])$ such that the following holds. If the observable $f$ is sufficiently smooth on $\mathbb{T}^{d}$ with $d\in\mathbb{N}^{+}\cup\{+\infty\}$, then for almost all rotations $\rho\in\mathbb{R}^{d}$, the uniform convergence of the corresponding§§§The normalization constant $A^{p,q}_{N}$ also changes accordingly. weighted Birkhoff average (1.2) exhibits a nearly linear-exponential rate. More precisely, the error is $\mathcal{O}\left({\exp({-c{N/\log N}})}\right)$ for some universal constant $c>0$. An explicit example of $w^{*}(x)$ is given by
@@
Taking into account (I)–(3) in Theorem 1.2, the conclusion of Theorem 1.3 is, in fact, not counterintuitive. The proof of this theorem essentially relies on the novel techniques established in [TL25a, Lemma 4.1], and the overall proof is similar to the quantitative results presented in [TL25a, Theorem 6.1]. However, it should be emphasized that no other existing techniques can effectively prove Theorem 1.3 beyond this approach. We have decided to present the complete proof in a separate paper and will simultaneously investigate the lower bounds of the convergence rates for the corresponding weighted Birkhoff averages.
@@ 1.1.2 Physically related problems revisited
    Computing rotations [DSSY17, DY18, SY18, SM20, MS21, BC23, DM23, BM24, CCGd24, MS25, SM25a, SM25b]: This is likely the most extensively utilized application of this weighted quasi-Monte Carlo method currently. The calculation of rotations–which are always well-defined and independent of the initial value–is often not an isolated issue. Instead, it may also serve as an essential intermediate component in other computational processes. High-precision rotations calculated through exponential acceleration enable one to often rely on experience to determine their arithmetic properties, thereby judging the corresponding dynamic properties. In the unweighted case, the convergence rate for computing rotations using standard Birkhoff averages is typically at most $\mathcal{O}\left(1/N\right)$, which stands in stark contrast to the exponential convergence achieved with weighting. The extensive use of this weighted quasi-Monte Carlo method in numerous studies strongly attests to its effectiveness.
@@
    Computing Lyapunov exponents [DSSY17, SM20, DM23, MS25, SM25a]: In dynamical systems, positive Lyapunov exponents, which measure average exponential stretching along orbits, typically characterize chaos. In contrast, for quasi-periodic/almost periodic systems, the Lyapunov exponents are zero. If all Lyapunov exponents are strictly less than zero, the system is certainly not chaotic but rather asymptotically stable. However, the standard orthogonalization-based method for computing Lyapunov exponents converges very slowly–typically at most $\mathcal{O}\left((\log N)/N\right)$–and may not converge at all. This weighted quasi-Monte Carlo method, while not guaranteed, can sometimes significantly accelerate the computation of Lyapunov exponents. However, there is currently no rigorous proof for this phenomenon, possibly due to singularities from the logarithmic term. This is an important issue we will need to address in the future.
@@
    Distinguishing between regular and chaotic orbits [SM20, MS21, DM23, MS25, SM25b]: This is one of the most fundamental questions in the field of dynamical systemsProf. James D. Meiss’s research group has made many fundamental contributions to (III) and (IV).. Although the Lyapunov exponent can be used to determine whether a system is chaotic to some extent, it is not a necessity. There are many other techniques that can achieve this goal, and most of them do not require the computation of specific Lyapunov exponents. This weighted quasi-Monte Carlo method can sometimes effectively distinguish between (weak/strong) chaotic orbits, island chains, and quasi-periodic/almost periodic orbits. For chaotic cases, in general, its convergence rate is no faster than that of the unweighted average of a random signal, with an error of $\mathcal{O}\left(N^{-1/2}\right)$. In contrast, for other regular cases, the convergence rate is typically exponential, provided the system is analytic, as detailed in Theorem 1.2.
@@
  • (VI)
In addition to the aforementioned applications, this weighted quasi-Monte Carlo method can also be employed to numerically validate the error of the “label functions” [RB25]. It has also been extended to other fields, including the study of numerical methods and other consistent approximations of the Euler system in gas dynamics, via a new concept of (S)-convergence [Fei21]. We would also like to mention [RB24], which employs reduced rank extrapolation to explore more general weighting methods (not necessarily in the form of (1.2), but depending on the specific systems), and achieves arbitrary exponential convergence similar to (I)–(3) in Theorem 1.2. In a study related to that of [RB24] but developed independently, [Sal24] introduced graph filters from the perspective of graph signal processing theory and explored the fast convergence in the ergodic theorem for Markov chains. We are convinced that this weighted quasi-Monte Carlo method holds potential applications far beyond what has been discussed in this paper. We look forward to seeing more works that delve into a broader range of problems related to dynamical systems to further demonstrate its effectiveness.

### v3

@@ 1.1.1 The weighted quasi-Monte Carlo method
$f=\sum\limits_{k\in\mathbb{Z}_{*}^{\infty}}{\widehat{f}(k)\exp\left({2\pi i\left\langle{k,x}\right\rangle}\right)}\quad\text{with}\quad\|f\|_{\sigma,\infty}:=\sum\limits_{k\in\mathbb{Z}_{*}^{\infty}}{\|{\widehat{f}(k)}\|_{\ell^{\infty}}{\exp\left(2\pi\sigma{{|k|_{\eta}}}\right)}}<+\infty.$
@@
    The discrete case: for any $0\neq k\in\mathbb{Z}_{*}^{\infty}$ and any $n\in\mathbb{Z}$, $\left|{\left\langle{k,\rho}\right\rangle-n}\right|>\gamma{\left({\prod\nolimits_{j\in{\mathbb{N}}}{(1+{{\left\langle j\right\rangle}^{\tau}}{{\left|{{k_{j}}}\right|}^{\tau}})}}\right)^{-1}}$, provided $\tau>1$;
@@
In contrast to the delicate results in finite dimensions, it seems that there is no known so-called “0–1 law” for nonresonance in infinite dimensions. However, the infinite-dimensional Diophantine nonresonance in Definition 1.4 is universal. More precisely, under the standard probability measure $\mathbb{P}$ induced by the product measure of the cube $\left[{1,2}\right]^{\mathbb{\infty}}$} (equivalent to $\mathbb{T}^{\infty}$}), the set of these rotations admits full probability measure [Bou05, MP21]. This is sufficient for us to establish the universal exponential convergence of weighted Birkhoff averages, via analytic observables. It should be emphasized that the universal nonresonance is not necessarily confined to the aforementioned Diophantine form. One can construct more general cases [TL24b], which we do not pursue here.
@@
      If the observable $f$ is “sub-super analytic” on $\mathbb{T}^{d}$, namely $\|\widehat{f}(k)\|_{\ell^{\infty}}=\mathcal{O}\left(\exp(-\|k\|_{\ell^{1}}^{v})\right)$ for any $v>1$, then for any given $0<\kappa<1$, when $p,q$ are sufficiently large, for almost all rotations $\rho\in\mathbb{R}^{d}$, the convergence can be of arbitrary exponential type. More precisely, the error is $\mathcal{O}\left({\exp({-{N^{\kappa}}})}\right)$.
@@
      If the observable $f$ is “super analytic” on $\mathbb{T}^{\infty}$, namely $\|\widehat{f}(k)\|_{\ell^{\infty}}=\mathcal{O}\left(\exp(-\exp(|k|_{\eta})\right)$, then for almost all rotations $\rho\in\mathbb{R}^{\infty}$, the convergence can be of exponential type. More precisely, the error is $\mathcal{O}\left({\exp({-{N^{\zeta_{4}}}})}\right)$ for any $0<{\zeta_{4}}<{\left({1+1/\min\left\{{p,q}\right\}}\right)^{-1}}$.
@@
      holds for some universal constant $c>0$, where ${\zeta_{5}}={\left({1+1/\min\left\{{p,q}\right\}}\right)^{-1}}$. Moreover, the universal and implied constants only depend on $T$ and ${\max_{0\leqslant n\leqslant T}}\left|{f\left(n\right)}\right|$§§§Here we regard $p$ and $q$ as fixed..
@@
As previously emphasized, although the results here are stated specifically for the case of toral translation, they can be readily formulated in a conjugate manner [DSSY17, DY18, DM23, RB24] to extend to general nonlinear dynamical systems (for example, those that can be smoothly conjugated to toral translations through KAM theory, etc.). Consequently, Theorem 1.2 has broad applicability, particularly in celestial mechanics, since the motion of celestial bodies is often quasi-periodic. For the sake of brevity, we prefer not to complicate the issue further.
@@
The terms “quasi-periodic” and “almost periodic” are used here because the corresponding rotations are nonresonant; for example, they satisfy certain finite-dimensional or infinite-dimensional Diophantine conditions as in Definitions 1.2 and 1.4. To simplify the exposition, we prefer not to state these concepts in Theorem 1.2. These rotations are universal in the sense of measure, meaning that the set of them admits full Lebesgue measure or full probability measure, as previously mentioned. This weighted quasi-Monte Carlo method is not only effective for quasi-periodic, almost periodic, and periodic cases, but also yields similar results for general decaying waves [TL25b, Theorems 1.1 and 1.2]. As a corollary, when the observational data can be decomposed into a linear combination of quasi-periodic, almost periodic, and periodic components, as well as decaying waves, this method also exhibits excellent acceleration effects. Here, we also mention that the “super-analyticity” imposed for the quasi-periodic case in (II)–(6) of Theorem 1.2 is not as strong as it might appear at first glance (of course, the corresponding “super-analyticity” in the finite-dimensional case is indeed strong). This is because, unlike the finite-dimensional case, the infinite-dimensional case requires consideration of the cardinality estimates brought by the $\mathbb{Z}_{*}^{\infty}$ lattice, which inevitably leads to stronger decay of individual Fourier coefficients.
@@
To facilitate the reader’s understanding of Theorem 1.2, we present the numerical simulations in Figure 4. For convenience, we denote the actual error in the $\ell^{\infty}$-norm as ${\rm{Error}}^{p,q}_{N}$. We consider a quasi-periodic orbit generated by $f(x)=\sin(2\pi x)+\cos(10\pi x)$, with an initial value $\theta=0$ and a rotation number $\rho=1/2\pi$, which is universal (i.e., not specifically constructed to achieve the “best” convergence). The choice of $f$ as a finite-order trigonometric polynomial is motivated by practical considerations: in numerical simulations, we cannot truly achieve infinity, and extremely high orders would lead to inaccurate computations. According to Theorem 1.2, specifically the part (I)–(4), the theoretical error is $\mathcal{O}\left(\exp(-cN^{\zeta_{2}})\right)$, where ${\zeta_{2}}={\left({1+1/\min\left\{{p,q}\right\}}\right)^{-1}}$ and $c>0$ is a universal constant that may depend on $p$ and $q$. To visualize the results more intuitively, we take the logarithm of both the actual error and the theoretical error (and thus the theoretical error becomes $\mathcal{O}\left(-N^{\zeta_{2}}\right)$, where ${\zeta_{2}}={\left({1+1/\min\left\{{p,q}\right\}}\right)^{-1}}$). This transformation results in curves that are more curved than $\mathcal{O}^{\#}(-N)$. We present four cases: $p=q=1/2$, $p=q=1$, $p=1$ and $q=2$, and $p=q=2$. It can be observed from Figure 4 that as $\min\{p,q\}$ increases, the exponential convergence rate of the actual error indeed accelerates as we proved, closely matching the changes in the order of the theoretical error (although we do not estimate the control constant $c>0$ as in [TL25b] for general decaying waves, which would have provided a more precise result). Notably, for the cases $p=q=1$ and $p=1$, $q=2$, the actual errors are almost identical, aligning well with our theoretical conclusions, as $\min\{1,1\}=\min\{1,2\}=1$ at this point.
In addition to the results in the general sense mentioned above, [TL24a, Theorems 2.1, 2.2, 3.1 and 3.2] and [TL24b, Theorems 3.1–3.4] also investigate results based on Liouville rotations. Specifically, for any Liouville type rotation, as long as the observable is sufficiently smooth, the weighted Birkhoff average can converge at an exponential rate. This is quite different from the phenomena observed in classical dynamical systems: Firstly, many dynamical systems essentially impose restrictions on the irrationality of rotations, such as the Brjuno condition, and do not allow arbitrary Liouville rotations. Secondly, even for the specific case of irrational translations, without weighting, such results are absolutely impossible according to the co-homological equation argument. However, when the regularity of the observable is insufficient, such weak nonresonance will reduce the convergence rate in the weighted case, which is entirely reasonable from the perspective of dynamical systems.
@@
There exist weighting functions $w^{*}(x)\in C_{0}^{\infty}([0,1])$ such that the following holds. If the observable $f$ is sufficiently smooth on $\mathbb{T}^{d}$ with $d\in\mathbb{N}^{+}\cup\{+\infty\}$, then for almost all rotations $\rho\in\mathbb{R}^{d}$, the uniform convergence of the corresponding***The normalization constant $A^{p,q}_{N}$ also changes accordingly. weighted Birkhoff average (1.2) exhibits a nearly linear-exponential rate. More precisely, the error is $\mathcal{O}\left({\exp({-c{N/\log N}})}\right)$ for some universal constant $c>0$. An explicit example of $w^{*}(x)$ is given by
@@
Taking into account (I)–(3) in Theorem 1.2, the conclusion of Theorem 1.3 is, in fact, not counterintuitive. The proof of this theorem essentially relies on the novel techniques established in [TL25b, Lemma 4.1], and the overall proof is similar to the quantitative results presented in [TL25b, Theorem 6.1]. However, it should be emphasized that no other existing techniques can effectively prove Theorem 1.3 beyond this approach. We have decided to present the complete proof in a separate paper and will simultaneously investigate the lower bounds of the convergence rates for the corresponding weighted Birkhoff averages.
@@ 1.1.2 Physically related problems revisited
    Computing rotations [DSSY17, DY18, SY18, SM20, MS21, BC23, DM23, BM24, CCGd24, BdJC25, BHS25, MS25, SM25a, SM25b]: This is likely the most extensively utilized application of this weighted quasi-Monte Carlo method currently. The calculation of rotations—which are always well-defined and independent of the initial value—is often not an isolated issue. Instead, it may also serve as an essential intermediate component in other computational processes. High-precision rotations calculated through exponential acceleration enable one to often rely on experience to determine their arithmetic properties, thereby judging the corresponding dynamic properties. In the unweighted case, the convergence rate for computing rotations using standard Birkhoff averages is typically at most $\mathcal{O}\left(N^{-1}\right)$, which stands in stark contrast to the exponential convergence achieved with weighting. The extensive use of this weighted quasi-Monte Carlo method in numerous studies strongly attests to its effectiveness.
@@
    Computing Lyapunov exponents [DSSY17, SM20, DM23, MS25, SM25a]: In dynamical systems, positive Lyapunov exponents, which measure average exponential stretching along orbits, typically characterize chaos. In contrast, for quasi-periodic/almost periodic systems, the Lyapunov exponents are zero. If all Lyapunov exponents are strictly less than zero, the system is certainly not chaotic but rather asymptotically stable. However, the standard orthogonalization-based method for computing Lyapunov exponents converges very slowly—typically at most $\mathcal{O}\left(N^{-1}\log N\right)$—and may not converge at all. This weighted quasi-Monte Carlo method, while not guaranteed, can sometimes significantly accelerate the computation of Lyapunov exponents. However, there is currently no rigorous proof for this phenomenon, possibly due to singularities from the logarithmic term. This is an important issue we will need to address in the future.
@@
    Distinguishing between regular and chaotic orbits [SM20, MS21, DM23, BHS25, BBC25, GZE25, MS25, SM25b, ZXH25]: This is one of the most fundamental questions in the field of dynamical systemsProf. James D. Meiss’s research group has made many fundamental contributions to (III) and (IV).. Although the Lyapunov exponent can be used to determine whether a system is chaotic to some extent, it is not a necessity. There are many other techniques that can achieve this goal, and most of them do not require the computation of specific Lyapunov exponents. This weighted quasi-Monte Carlo method can sometimes effectively distinguish between (weak/strong) chaotic orbits, island chains, and quasi-periodic/almost periodic orbits. For chaotic cases, in general, its convergence rate is no faster than that of the unweighted average of a random signal, with an error of $\mathcal{O}\left(N^{-1/2}\right)$. In contrast, for other regular cases, the convergence rate is typically exponential, provided the system is analytic, as detailed in Theorem 1.2.
@@
In addition to the aforementioned applications, this weighted quasi-Monte Carlo method can also be employed to numerically validate the error of the “label functions” [RB25]. It has also been extended to other fields, including the study of numerical methods and other consistent approximations of the Euler system in gas dynamics, via a new concept of (S)-convergence [Fei21]. We would also like to mention [RB24], which employs reduced rank extrapolation to explore more general weighting methods (not necessarily in the form of (1.2), but depending on the specific systems), and achieves arbitrary exponential convergence similar to (I)–(3) in Theorem 1.2. In a study related to that of [RB24] but developed independently, [Sal24] introduced graph filters from the perspective of graph signal processing theory and explored the fast convergence in the ergodic theorem for Markov chains. Recently, [BBC25] presented a comprehensive integration of this weighted quasi-Monte Carlo method with five data-driven algorithms from dynamical systems, accompanied by extensive examples. We are convinced that this weighted quasi-Monte Carlo method holds potential applications far beyond what has been discussed in this paper. We look forward to seeing more works that delve into a broader range of problems related to dynamical systems to further demonstrate its effectiveness.